OpenStudy (anonymous):

I have a questions that doesn't appear to be an equation. Not sure how to solve. What is the solution of sq rt.(5x+1) The possible sol'n are x =0; x=16 and x=0; x=16; x=16 and x=1

OpenStudy (klimenkov):

If you want to get a solution type an equation firstly.

OpenStudy (anonymous):

Shouldn't there be an equals sign somewhere in the initial equation of sqrt*x+) ?

OpenStudy (anonymous):

sorry sqrt(5x+1)

OpenStudy (anonymous):

there is no = sign just the sq rt sign with 5x+1 under it Can it not be solved unless there is an = sign. Maybe something is missing.

OpenStudy (anonymous):

Maybe look higher up sometimes my math textbook says like for all of these equations solve when equal to five or something like that.

OpenStudy (anonymous):

I don't get what u r saying

OpenStudy (anonymous):

ahhh I think I know what they mean now I think the equation is in the form y=sqrt(5x+1) which means they want you to solve for when it crosses the x axis. See if that sounds like something you have/are doing in class.

OpenStudy (anonymous):

my lesson is solving sq rt and other radical equations

OpenStudy (anonymous):

Maybe it means which solution sets are possible....

OpenStudy (anonymous):

Some constraint....

OpenStudy (anonymous):

Unfortunately I cannot help you as a) I do not fully understand what you are asking and b) I have to go I have class in 15 minutes

OpenStudy (anonymous):

I would like to see wording of exact question...

OpenStudy (anonymous):

exact wording is: What is the solution of sq. rt (5x+1)?

OpenStudy (anonymous):

The whole question....

OpenStudy (anonymous):

that is the whole question

OpenStudy (anonymous):

What about the solutions part?

OpenStudy (anonymous):

possbile answers are x=0; x=16 and x=0; x=16; x=16 and x=1

OpenStudy (phi):

It might be a typo in the question. The answers you gave imply sqrt(5x+1)= 1 or maybe sqrt(5x+1)= 9 are possible equations.

OpenStudy (anonymous):

What is the solution of sq. rt (5x+1)? Possbile answers are x=0; x=16 and x=0; x=16; x=16 and x=1 That's what it says?

OpenStudy (anonymous):

yea because if you solve for zeros you get 5x+1=0 5x=-1 x=-1/5

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

I tried solving for 0

OpenStudy (anonymous):

I was thinking it's a typo but wasn't sure. Thought maybe I was missing something

OpenStudy (anonymous):

As far as I can tell you seem to be missing a crucial piece of this question somewhere.

OpenStudy (phi):

I would bet your are a victim here. If you post a scan of the page, it might make it obvious to us...